### v1

# CTP-SCU/2025011

Abstract
We provide a field-theoretic method to calculate entanglement entropy of CFT in all dimensions, for arbitrary shape of entangling surfaces. Ryu-Takayanagi formula is manifested in this formalism.
preprint: preprintnumbers
I Introduction
Entanglement properties of quantum systems are playing an increasingly crucial role in physics. An intrinsic quantity counting the degrees of freedom of entanglement is the von Neumann entropy, also known as entanglement entropy. It is defined by $S_{A}=-\mathrm{Tr}\,\rho_{A}\log\rho_{A}$ for a reduced density matrix $\rho_{A}$ associated with a subsystem $A$, where the total system is bipartitioned into two subsystems, $A$ and its complement $A^{c}$. In quantum field theory (QFT), one usually considers the entropy of entanglement between two adjacent and complementary subregions $A$ and $A^{c}$, denoted as $S(A:A^{c})$. Although the calculation of entanglement entropy in two dimensional CFT has been well established using the replica trick [1, 2], no approach currently exists for performing such calculations in higher dimensional CFTD for $D>2$.
The entropy $S(A:A^{c})$ typically suffers an ultraviolet (UV) divergence due to the very intense entanglement between contiguous fields. This divergent behavior renders it extremely challenging, if not entirely impossible, to derive exact relations between QFT and gravity. A natural question arises: are there any finite elementary entanglement entropies? For such quantities to exist, two conditions should be satisfied:
  • The entangled regions must be disjoint.
  • The system has to be a pure state to allow the calculation of the entanglement entropy.
It is not hard to understand that the annular CFT2 in Figure 1 precisely meets these two conditions. In prior work with collaborators [3, 4, 5], we demonstrated that the entanglement entropy $S(A:B)$ between disjoint intervals $A$ and $B$ is indeed finite, for static, covariant and thermal configurations, respectively. Notably, $S(A:B)$ can be equivalently interpreted as the entanglement entropy of the mixed state $\rho_{AB}$ in the infinite system, as illustrated in Figure 2. The well-known divergent entanglement entropies for adjacent intervals emerge naturally as the adjacent limits of disjoint $S(A:B)$.
Figure 1: The Euclidean pure state density matrix for the annular CFT2. Two distinct intervals $A$ and $B$ have finite entanglement entropy $S(A:B)$.
Figure 2: After subtracting segments $C$ and $D$ with two discs in the infinite system, we obtain an annular region in which $A$ and $B$ are in a pure entangled state $\psi_{AB}$. It is identical to the annular CFT2.
A critical insight is the straightforward generalization of the annular CFT2 to higher dimensions: CFTD on a solid torus $\mathbb{B}^{D-1}\times S^{1}$, see Figure 3, where $\mathbb{B}^{D-1}$ is a $(D-1)$-dimensional ball. Since the replica of a solid torus is still a solid torus, calculating entanglement entropies simplifies to computing the partition function on this solid torus! The usual divergent entanglement entropies for adjacent regions are again easily obtained by taking adjacent limits. In this paper, we only study the ball shaped entangling region. But one can easily see that this framework works for entangling surfaces of arbitrary shape.
For $D>2$, the relationship between the solid torus CFTD and the infinite CFTD on $\mathbb{R}^{D}$ mirrors the $D=2$ case. Consider calculating the entanglement entropy between two codimension-one disjoint discs $A$ and $B$ in $\mathbb{R}^{D}$, corresponding to the mixed state $\rho_{AB}$ of the infinite system. Analogous to the two-dimensional scenario, removing a D-dimensional solid torus between $A$ and $B$ yields also a D-dimensional solid torus depicted in Figure 3 — precisely the SUBTRACTION procedure introduced in [3].
The remainder of this paper is outlined as follows. In section II, we derive the explicit expression of the entanglement entropy between two disjoint balls in all dimensions. In section III, by taking the adjacent limits, the divergent entanglement entropies for adjacent balls in all dimensions are obtained, where the area law [6, 7] is manifested. In section IV, we provide examples in CFT2 and CFT4. In section V, we show how RT formula [8, 9] naturally emerges in this framework and discuss arbitrary shape of entangling surfaces. In section VI, we find perfect agreement of our results with the holographic calculation.
II Entanglement entropy between two disjoint balls
For a CFT in $D$-dimensional Euclidean spacetime with the metric $\mathrm{d}s_{0}^{2}=\mathrm{d}t_{\text{E}}^{2}+\mathrm{d}y^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2}$, we consider quantum fields living in a bounded region,
:=$\mathcal{B}:=\left\{(t_{\text{E}},y,x)|\left(\sqrt{t_{\text{E}}^{2}+y^{2}}-\frac{l_{2}+l_{1}}{2}\right)^{2}+\sum_{i=1}^{D-2}x_{i}^{2}<\left(\frac{l_{2}-l_{1}}{2}\right)^{2}\right\},$ (1)
with $l_{2}>l_{1}>0$. At the time slice $t_{\text{E}}=0$, $\mathcal{B}$ consists of two disjoint $(D-1)$-balls $A$ and $B$ with the same radius, as shown in Figure 3. The bounded region $\mathcal{B}$ is intuitively a solid torus $\mathbb{B}^{D-1}\times S^{1}$, where $\mathbb{B}^{D-1}$ is a $(D-1)$-dimensional ball. Up to a Weyl transformation $\mathrm{d}s_{0}^{2}\rightarrow\mathrm{d}s_{0}^{2}/r^{2}$, the spacetime metric becomes
$\mathrm{d}s^{2}=\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\right)/r^{2},$ (2)
in terms of coordinates transformations, $t_{\text{E}}=r\sin\theta$, $y=r\cos\theta$. In this parameterization, we can treat $\theta$ as imaginary time with a periodicity $2\pi$, and the CFT living in $\mathcal{B}$ can be thought as a thermal field theory with the inverse temperature $\beta=2\pi$. The partition function of the CFT can be formally written as a trace over the Hilbert space:
$Z=\mathrm{Tr}\,e^{-2\pi H}.$ (3)
The Hamiltonian operator $H$ conjugate to the imaginary time $\theta$ is defined by
$H=\int_{\mathbb{B}^{D-1}}T^{\theta\theta}$
where $T$ is the stress-energy tensor.
Figure 3: Entangling regions $A$ and $B$ (colored regions) in the bounded region $\mathcal{B}$ (blue region).
Starting point of our computations is the observation that the vacuum entanglement entropy between $A$ and $B$ can be computed by the replica trick [1]. Such techniques allow us to write the Rényi entropy as
$S^{(n)}=\frac{1}{1-n}\log\left[\frac{Z_{\mathcal{B}_{n}}}{Z^{n}}\right],$ (4)
where $\mathcal{B}_{n}$ is made by cyclically gluing $n$ copies of $\mathcal{B}$ with a cut along $A$ (or equivalently $B$), and the metric of $\mathcal{B}_{n}$ might be given by
$\mathrm{d}s_{(n)}^{2}=n^{2}\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\right)/r^{2}.$ (5)
The partition function of CFT living on $\mathcal{B}_{n}$ is then
$Z_{\mathcal{B}_{n}}=\mathrm{Tr}\,e^{-2n\pi H_{(n)}},$ (6)
with the corresponding Hamiltonian operator $H_{(n)}$. We will focus on the partition function that is dominated by the ground state,
$Z\simeq e^{-2\pi\langle H\rangle},\quad Z_{\mathcal{B}_{n}}\simeq e^{-2n\pi\langle H_{(n)}\rangle}.$ (7)
$\langle H\rangle$ denotes the vacuum expectation value of $H$. Note that $\mathcal{B}$ is related to the replicated manifold $\mathcal{B}_{n}$ by a rescaling of imaginary time $\theta\rightarrow n\theta$, which implies that [^1] Note the stress-energy tensor is covariant under coordinates transformations, $T^{\theta^{\prime}\theta^{\prime}}\partial_{\theta^{\prime}}\partial_{\theta^{\prime}}=T^{\theta\theta}\partial_{\theta}\partial_{\theta}$, and $\theta^{\prime}=n\theta$.
$\langle H_{(n)}\rangle=\frac{\langle H\rangle}{n^{2}}.$ (8)
Then, the Rényi entropy becomes
$S^{(n)}=\frac{1}{1-n}\log\left[\frac{e^{-2\pi\langle H\rangle/n}}{e^{-2n\pi\langle H\rangle}}\right]=-2\pi\left(1+\frac{1}{n}\right)\langle H\rangle.$ (9)
The entropy $S(A:B)$ of entanglement between two disjoint balls $A,B$ therefore reads
$S(A:B)=S^{(1)}=-4\pi\langle H\rangle.$ (10)
We stress that the Hamiltonian $H$ is conjugate to $\theta$ but not to $t_{\text{E}}$, so it is actually dimensionless.
The entanglement entropy $S(A:B)$ simply depends on the ground state energy $\langle H\rangle$ localized in a spatial ball $\mathbb{B}^{D-1}$:
$\langle H\rangle=\int_{\mathbb{B}^{D-1}}\mathcal{E}_{\text{vac}},$ (11)
where $\mathcal{E}_{\text{vac}}$ is the Casimir energy density that is usually a negative constant in our discussions. Consequently, the entanglement entropy $S(A:B)$ is given by
$S(A:B)$ $=$ $-4\pi\mathcal{E}_{\text{vac}}\mathrm{Vol}(\mathbb{B}^{D-1}),$ (12)
$=$ $-4\mathcal{E}_{\text{vac}}\pi^{\frac{D}{2}}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{l_{2}}{l_{1}}-1\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;1-\frac{l_{2}}{l_{1}}\right).$
with the standard hypergeometric function $\,{}_{2}F_{1}$. The parameters $l_{1}$ and $l_{2}$ are the inner and outer radii of the solid torus, as indicated in the left panel of Figure 5.
Figure 4: Entangling regions (colored regions) are bounded by ($D-2$)-spheres (black circles), respectively. Left panel (the cavity configuration): A ($D-1$)-ball living in a spheric cavity is entangled with the region outside the cavity. Right panel (the juxtaposed configuration): Two disjoint ($D-1$)-balls with different radii are entangled.
Generally, there exist two different configurations of our interest, namely, the cavity configuration and the juxtaposed configuration illustrated in Figure 4. Each configuration is uniquely fixed by boundaries of $A,B$, i.e., two disjoint ($D-2$)-spheres. Every ($D-1$)-ball is bounded by a ($D-2$)-sphere $\Sigma(\vec{x},r)$ located at $\vec{x}\in\mathbb{R}^{D-1}$ with the radius $r$, respectively. An elegant substitute for conformality is the inversive product $\varrho$ of two spheres $\Sigma(\vec{x},r)$ and $\Sigma(\vec{x}^{\prime},r^{\prime})$ [10]
$\varrho=\left|\frac{r^{2}+r^{\prime 2}-|\vec{x}-\vec{x}^{\prime}|^{2}}{2rr^{\prime}}\right|.$ (13)
In the case that two disjoint ($D-1$)-balls $A,B$ share the same radius, we have
$\frac{l_{2}}{l_{1}}=\frac{\sqrt{\varrho+1}+\sqrt{2}}{\sqrt{\varrho+1}-\sqrt{2}}.$ (14)
The inversive product $\varrho$ is a conformal invariant, as it remains invariant under global conformal transformations. If two configurations have different inversive products, they cannot be conformally mapped onto each other. The set of all configurations falls into classes of conformally equivalent regions, with each class characterized by the inversive product $\varrho$ of that class. Consequently, any two configurations sharing the same $\varrho$ must be conformally equivalent. To compute the entanglement entropy $S(A:B)$ for any two disjoint spheres in general configurations, it suffices to evaluate the inversive product $\varrho$ of the spheres. Thus, we conclude that the entanglement entropy between any two disjoint spheres is given by [^2] For the time dependent covariant scenario, simply replace $\vec{x}$ with $x^{\mu}$.
$S=-4\mathcal{E}_{\text{vac}}\pi^{\frac{D}{2}}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{2\sqrt{2}}{\sqrt{\varrho+1}-\sqrt{2}}\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;\frac{2\sqrt{2}}{\sqrt{2}-\sqrt{\varrho+1}}\right).$ (15)
III Emergent area law in various dimensions
In this section, we take the adjacent limit of the general finite entanglement entropy (15) of disjoint balls to give the usual divergent entanglement entropies of adjacent balls.
Figure 5: Left panel: The time zero slice of $\mathcal{B}$ in $D$ dimensions, where $A$ and $B$ are two disjoint ($D-1$)-balls with the same radius (colored regions) living at $t_{\text{E}}=0$. Right panel: The ($D-1$)-ball $A$ with the radius $R-\epsilon$ is in a spheric cavity with the radius $R+2\epsilon$, and the region $B$ is outside the cavity. Here, $\epsilon$ plays a role of the UV cutoff.
We first consider the entanglement entropy $S(A:B)$ for two disjoint ($D-1$)-balls $A,B$ with the same radius in the juxtaposed configuration, as shown in Figure 5. It would be convenient to write down $S(A:B)$ in various dimensions:
$D=2:$ $4\pi\mathcal{E}_{2}\log\frac{l_{2}}{l_{1}},$
$D=3:$ $4\pi^{2}\mathcal{E}_{3}\left(\sqrt{\frac{l_{2}}{l_{1}}}+\sqrt{\frac{l_{1}}{l_{2}}}\right)-8\pi^{2}\mathcal{E}_{3},$
$D=4:$ $2\pi^{2}\mathcal{E}_{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-4\pi^{2}\mathcal{E}_{4}\log\frac{l_{2}}{l_{1}},$
$D=5:$ $\frac{\pi^{3}\mathcal{E}_{5}}{3}\left[\left(\frac{l_{2}}{l_{1}}\right)^{3/2}+\left(\frac{l_{1}}{l_{2}}\right)^{3/2}-9\left(\sqrt{\frac{l_{2}}{l_{1}}}+\sqrt{\frac{l_{1}}{l_{2}}}\right)\right]+\frac{16\pi^{3}\mathcal{E}_{5}}{3},$
$D=6:$ $\frac{\pi^{3}\mathcal{E}_{6}}{6}\left[\left(\frac{l_{2}}{l_{1}}\right)^{2}-\left(\frac{l_{1}}{l_{2}}\right)^{2}-8\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)\right]+2\pi^{3}\mathcal{E}_{6}\log\frac{l_{2}}{l_{1}},$ (16)
where we denote $\mathcal{E}_{\text{vac}}=-\mathcal{E}_{D}<0$ and remember that $\mathcal{E}_{D}$ is actually different in various dimensions.
Then, we return to the general $D$-dimensional cavity configuration depicted in the right panel of Figure 5, where $A$ is almost adjacent to $B$, more precisely implying that
$\vec{x}=\vec{x}^{\prime},\quad r=R+\epsilon,\quad r^{\prime}=R-\epsilon.$ (17)
In this setup, the inversive product becomes
$\varrho=\frac{R^{2}+\epsilon^{2}}{R^{2}-\epsilon^{2}}.$ (18)
Plugging $\varrho$ into Equation (16), we obtain the corresponding entanglement entropy in various dimensions:
$D=2:$ $8\pi\mathcal{E}_{2}\log\frac{2R}{\epsilon}+\mathcal{O}(\epsilon),$
$D=3:$ $8\pi^{2}\mathcal{E}_{3}\left(\frac{R}{\epsilon}-1\right),$
$D=4:$ $2\pi^{2}\mathcal{E}_{4}\left(\frac{4R^{2}}{\epsilon^{2}}-2-4\log\frac{2R}{\epsilon}\right)+\mathcal{O}\left(\epsilon\right),$
$D=5:$ $\frac{8\pi^{3}\mathcal{E}_{5}}{3}\left(\frac{R^{3}}{\epsilon^{3}}-\frac{3R}{\epsilon}+2\right),$
$D=6:$ $\frac{\pi^{3}\mathcal{E}_{6}}{3}\left(\frac{8R^{4}}{\epsilon^{4}}-\frac{24R^{2}}{\epsilon^{2}}+9+12\log\frac{2R}{\epsilon}\right)+\mathcal{O}\left(\epsilon\right).$ (19)
In $D=2$, the result is the well-known $\log$-law for entanglement entropy. In $D\geq 3$, we can immediately observe the area law of entanglement entropy, which arises as a special limit of our general consideration. We emphasize that in even dimensions, the finite term is physically irrelevant since they can be absorbed into the UV cutoff $\epsilon$. But in odd dimensions, finite terms of entanglement entropies become scheme-independent and play a crucial role in the F-theorem [11, 12, 13]. This is manifested by the exactness of our expressions for entanglement entropy in odd dimensions.
IV Examples in CFT2 and CFT4
We provide two consistency checks in this section. Consider a CFT4 living in the 4-dimensional conformally flat spacetime with the metric $\mathrm{d}s^{2}=\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{2}\mathrm{d}x_{i}^{2}\right)/r^{2}$, where the vacuum expectation value of the stress tensor has been explicitly calculated in [14],
$\mathcal{E}_{\text{vac}}=\left\langle T^{\theta\theta}\right\rangle=-\frac{3a+c}{8\pi^{2}},$ (20)
where $a$ and $c$ are different coefficients of the trace anomaly [^3] In our notations, the trace anomaly is $\left\langle T_{\;\mu}^{\mu}\right\rangle=\frac{1}{(4\pi)^{2}}\left[c(C_{\mu\nu\rho\sigma}^{2}+\frac{2}{3}\nabla^{2}\mathcal{R})-a\mathcal{G}\right]$, with the Weyl tensor $C_{\mu\nu\rho\sigma}$, the Ricci scalar $\mathcal{R}$ and the topological Euler density $\mathcal{G}.$. Plugging Equation (20) and $D=4$ into Equation (12), the entanglement entropy between $A$ and $B$ is given by
$S(A:B)=\frac{3a+c}{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-\frac{3a+c}{2}\log\frac{l_{2}}{l_{1}}.$ (21)
Note that $A$, $B$ are two disjoint 3-balls with the same radius. In the 4-dimensional $\mathcal{N}=4$ $SU(N)$ super Yang-Mills gauge theory, the anomaly coefficients are [15, 16]
$a=\frac{N^{2}}{4},\quad c=\frac{N^{2}}{4}.$ (22)
Hence
$S(A:B)=\frac{N^{2}}{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-\frac{N^{2}}{2}\log\frac{l_{2}}{l_{1}}.$ (23)
Traditionally, that $A$ and $B$ are two adjacent and complementary regions is of our main interest, as depicted in the right panel of Figure 5. This case is conformally equivalent to two disjoint 3-balls with the same radius, due to the fact that any global conformal transformations are sphere-preserving [10], which has been analyzed in a more general way in Section III. Consequently, in Figure 5, the left panel and the right panel are conformally equivalent by setting that
$\frac{l_{2}}{l_{1}}\simeq\left(\frac{2R}{\epsilon}\right)^{2}-2+\mathcal{O}(\epsilon),$ (24)
if we take the adjacent limit $\epsilon\rightarrow 0$. Such a setup finally gives us the area law of entanglement entropy:
$S(A:A^{c})\simeq N^{2}\left[\left(\frac{R}{\epsilon}\right)^{2}-\log\frac{2R}{\epsilon}-2\right]+\mathcal{O}(\epsilon).$ (25)
This is actually entanglement entropy of a ball with the radius $R$, which precisely matches the holographic prediction given in [8].
As $D=2$, the metric reduces to $\mathrm{d}s^{2}=\mathrm{d}\theta^{2}+\mathrm{d}r^{2}/r^{2}$, which is just a flat metric on the Euclidean cylinder. Now, $A=\{(\theta,r)|l_{1}<r<l_{2},\theta=\pi\}$ and $B=\{(\theta,r)|l_{1}<r<l_{2},\theta=0\}$ are two disjoint intervals. The Casimir energy density is related to the central charge [17]:
$\mathcal{E}_{\text{vac}}=-\frac{c}{24\pi}.$ (26)
The CFT2 result presented in our previous works [3, 4] is immediately reproduced,
$S(A:B)=\frac{c}{6}\log\frac{l_{2}}{l_{1}}.$ (27)
Its adjacent limit gives the entanglement entropy of a single interval with the length $2R$,
$S(A:A^{c})=\frac{c}{3}\log\frac{2R}{\epsilon}+\mathcal{O}(\epsilon).$
V Shape independence and Ryu-Takayanagi formula
We present two important indications or extensions in this section.
Shape independence
Even though we only study the spheric entangling surface in this paper, our approach is clearly independent of the specific shape of entangling regions $A$ and $B$. In other words, one can replace $A$ and $B$ with generally shaped regions, denoted as $\mathbb{X}$. The relevant entanglement entropy would then be
$S(A:B)$ $=$ $-4\pi\mathcal{E}_{\text{vac}}\mathrm{Vol}(\mathbb{X}).$ (28)
However, it should be noted that to apply this formula directly, $A$ and $B$ must be reflection-symmetric with respect to each other, as only under this condition the thermal partition function can be formulated.
To calculate the entanglement entropy between two similar, disjoint regions (same shape, different sizes) with arbitrary positions, one needs to map the configuration to the reflection-symmetric one with some conformal map. For any shape of the entangling region, there should exist a set of inversive products analogous to (13).
For adjacent scenarios, it is easy to understand that, by choosing some simple setup and then taking adjacent limit, the usual divergent entanglement entropy for arbitrarily shaped region can be obtained straightforwardly.
Manifestation of Ryu-Takayanagi formula
Since any pair of similar, disjoint regions falls into the reflection-symmetric case, the entanglement entropy is proportional to the area $\mathrm{Vol}(\mathbb{X})$ defined within the hyperbolic metric $\mathrm{d}s^{2}=(\mathrm{d}r^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2})/r^{2}.$ Note that the subregion $\mathbb{X}$ could always be understood as a ($D-1$)-dimensional minimal surface embedded in higher-dimensional AdSD+1 space. Therefore, the RT formula saying that the entanglement entropy of CFTD is proportional to the area of the codimension-2 minimal surface in the AdSD+1 spacetime has been manifested in current method.
VI Holographic entanglement entropy
Figure 6: At the time slice $t_{\text{E}}=0$, two subregions $A,B$ are living on the AdS boundary $z=0$, while the minimal cross-section $E_{W}$ (gray region) of the entanglement wedge (the region bounded by the blue tube) of $A$ and $B$ is in the AdS bulk.
It would be nice to explore the gravity dual of finite entanglement entropy in the context of AdS/CFT correspondence [18, 19, 20]. Recall that a CFT in $D$-dimensional Euclidean spacetime with the metric $\mathrm{d}s_{0}^{2}=\mathrm{d}t_{\text{E}}^{2}+\mathrm{d}y^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2}$ is dual to a ($D+1$)-dimensional Euclidean AdS spacetime with the metric
$\mathrm{d}s^{2}=\frac{\ell_{\text{AdS}}^{2}}{z^{2}}\left(\mathrm{d}z^{2}+\mathrm{d}t_{\text{E}}^{2}+\mathrm{d}y^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\right).$ (29)
The entanglement entropy between two disjoint $(D-1)$-balls $A$ and $B$ with the same radius is dual to a ball-shaped region at the time slice $t_{\text{E}}=0$ of the AdSD+1 bulk, i.e., the entanglement wedge cross-section,
$E_{W}:=\left\{(x,z)|\left(z-\frac{l_{2}+l_{1}}{2}\right)^{2}+\sum_{i=1}^{D-2}x_{i}^{2}<\left(\frac{l_{2}-l_{1}}{2}\right)^{2}\right\}.$ (30)
The area of the entanglement wedge cross-section is given by
$\mathsf{A}=\ell_{\text{AdS}}^{D-1}\pi^{\frac{D}{2}-1}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{l_{2}}{l_{1}}-1\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;1-\frac{l_{2}}{l_{1}}\right).$ (31)
According to the RT formula, the corresponding entanglement entropy is
$S(A:B)=\frac{\mathsf{A}}{4G_{N}^{(D+1)}}=\frac{\ell_{\text{AdS}}^{D-1}}{4G_{N}^{(D+1)}}\pi^{\frac{D}{2}-1}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{l_{2}}{l_{1}}-1\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;1-\frac{l_{2}}{l_{1}}\right),$ (32)
with the ($D+1$)-dimensional gravitational constant $G_{N}^{(D+1)}$. This coincides with our field-theoretic result (15). For instance, in $D=4$, one has
$S(A:B)$ $=$ $\frac{3a+c}{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-\frac{3a+c}{2}\log\frac{l_{2}}{l_{1}}$ (33)
$=$ $\frac{\pi\ell_{\text{AdS}}^{3}}{8G_{N}^{(5)}}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}-2\log\frac{l_{2}}{l_{1}}\right).$ (34)
with two central charges $a=c=\pi\ell_{\text{AdS}}^{3}/8G_{N}^{(5)}$ [21, 22].
It is straightforward to extend the entanglement wedge cross-section to the general case, such as the cavity configuration we have addressed in the CFT side. This is because the AdS bulk and the dual CFT always share the same conformal symmetry. The entanglement wedge cross-section in the general configuration can be conformally mapped onto the reflection-symmetric case studied above.
When taking the adjacent limit, one immediately reproduces the original RT formula for the adjacent scenario with divergent entanglement entropy.
VII Conclusion
In summary, we introduced a general framework to compute the entanglement entropy in CFTD for $D\geq 2$. We provided an explicit expression for the entanglement entropy of ball-shaped entangling regions in CFTD. This method is applicable to entangling regions of arbitrary shape. Remarkably, this method is substantially simpler than the traditional one used to calculate CFT2 entanglement entropy. The RT formula emerges naturally within this formalism.
In CFT2, the divergent entanglement entropy between two adjacent and complementary subsystems has been extensively studied. However, the finite entanglement entropy between two disjoint regions has remained underexplored until recently. From our perspective, the latter is of greater fundamental importance: it provides a pair of exactly equal dual quantities in the correspondence between gravity and CFT, whereas the divergent adjacent configuration represents merely a special limiting case of this disjoint configuration.
Acknowledgements.
Acknowledgments.
This work is supported by NSFC (Grant No. 12275184).

### v2

@@
 Abstract
-We provide a field-theoretic method to calculate entanglement entropy of CFT in all dimensions, for arbitrary shape of entangling surfaces. Ryu-Takayanagi formula is manifested in this formalism.
+We provide a field-theoretic method to calculate entanglement entropy of CFT in all dimensions.
+This method works for entangling surfaces of arbitrary shape.
+The formalism manifests a field-theoretic proof of the Ryu-Takayanagi formula.
 preprint: preprintnumbers
@@
 It is defined by $S_{A}=-\mathrm{Tr}\,\rho_{A}\log\rho_{A}$ for a reduced density matrix $\rho_{A}$ associated with a subsystem $A$, where the total system is bipartitioned into two subsystems, $A$ and its complement $A^{c}$.
-In quantum field theory (QFT), one usually considers the entropy of entanglement between two adjacent and complementary subregions $A$ and $A^{c}$, denoted as $S(A:A^{c})$.
-Although the calculation of entanglement entropy in two dimensional CFT has been well established using the replica trick [1, 2], no approach currently exists for performing such calculations in higher dimensional CFTD for $D>2$.
-The entropy $S(A:A^{c})$ typically suffers an ultraviolet (UV) divergence due to the very intense entanglement between contiguous fields.
+In QFT, one usually considers the entanglement entropy between two adjacent and complementary subregions $A$ and $A^{c}$, denoted as $S_{\text{adj}}(A:A^{c})$, where the subscript “adj” stands for adjacent entangling regions.
+Although the calculation of entanglement entropy in two dimensional CFT has been well established using the replica trick [1, 2], no general approach currently exists for performing such calculations in higher dimensional CFTD for $D>2$.
+The entropy $S_{\text{adj}}(A:A^{c})$ typically suffers an ultraviolet (UV) divergence due to the very intense entanglement between contiguous fields.
 This divergent behavior renders it extremely challenging, if not entirely impossible, to derive exact relations between QFT and gravity.
-A natural question arises: are there any finite elementary entanglement entropies?
-For such quantities to exist, two conditions should be satisfied:
+However, for a UV-complete theory such as a CFT, all physical quantities must possess well-defined UV-finite forms.
+Divergences and regulators, rather than being intrinsic components of the theory itself, should only emerge as artifacts of specific limiting procedures.
+A natural question thus arises: Do there exist any finite elementary entanglement entropies in CFT?
+Here, the term elementary is used in contrast to compound quantities such as mutual information, relative entropy, and the like — which are constructed from combinations of elementary entanglement measures.
+For such finite elementary entanglement entropies to exist, two conditions should be satisfied:
   • The entangled regions must be disjoint.
   • The system has to be a pure state to allow the calculation of the entanglement entropy.
-It is not hard to understand that the annular CFT2 in Figure 1 precisely meets these two conditions.
-In prior work with collaborators [3, 4, 5], we demonstrated that the entanglement entropy $S(A:B)$ between disjoint intervals $A$ and $B$ is indeed finite, for static, covariant and thermal configurations, respectively.
-Notably, $S(A:B)$ can be equivalently interpreted as the entanglement entropy of the mixed state $\rho_{AB}$ in the infinite system, as illustrated in Figure 2.
-The well-known divergent entanglement entropies for adjacent intervals emerge naturally as the adjacent limits of disjoint $S(A:B)$.
-Figure 1: The Euclidean pure state density matrix for the annular CFT2. Two distinct intervals $A$ and $B$ have finite entanglement entropy $S(A:B)$.
-Figure 2: After subtracting segments $C$ and $D$ with two discs in the infinite system, we obtain an annular region in which $A$ and $B$ are in a pure entangled state $\psi_{AB}$.
-It is identical to the annular CFT2.
-A critical insight is the straightforward generalization of the annular CFT2 to higher dimensions: CFTD on a solid torus $\mathbb{B}^{D-1}\times S^{1}$, see Figure 3, where $\mathbb{B}^{D-1}$ is a $(D-1)$-dimensional ball.
+The annular CFT2 in Figure 1 precisely meets these two conditions.
+Note the time direction is upward.
+The lower half annulus corresponds to $|\psi\rangle$, and the upper half annulus corresponds to $\langle\psi|$.
+The density matrix of this pure state is $\rho=|\psi\rangle\langle\psi|$.
+In prior collaborations [3, 4, 5], we demonstrated that the entanglement entropy $S_{\text{disj}}(A:B)$ between disjoint intervals $A$ and $B$ in this annular CFT is indeed finite, for static, covariant and thermal configurations, respectively.
+The subscript “disj” denotes disjoint entangling regions.
+Notably, referring to Figure 2, as shown in [3, 4], $S_{\text{disj}}(A:B)$ can be equivalently interpreted as the entanglement entropy between two disjoint intervals in the infinite system.
+The well-known divergent entanglement entropies $S_{\text{adj}}(A:B)$ for adjacent intervals emerge naturally as the adjacent limits of disjoint $S_{\text{disj}}(A:B)$.
+Figure 1: The Euclidean pure state density matrix $\rho=|\psi\rangle\langle\psi|$ for the annular CFT2. Two distinct intervals $A$ and $B$ have finite entanglement entropy $S(A:B)$.
+Figure 2: As shown in [3, 4], after subtracting segments $C$ and $D$ with two discs in the infinite system, we obtain an annular region in which AA and $B$ are in a pure entangled state $\psi_{AB}$.
+It is identical to the annular CFT2 in Figure 1.
+A critical insight is the straightforward generalization of the annular CFT2 to higher dimensions: CFTD on a solid torus $\mathbb{B}^{D-1}\times S^{1}$, see Figure 3, where $\mathbb{B}^{D-1}$ is a (D1)(D-1)-dimensional ball.
 Since the replica of a solid torus is still a solid torus, calculating entanglement entropies simplifies to computing the partition function on this solid torus!
@@
 For $D>2$, the relationship between the solid torus CFTD and the infinite CFTD on $\mathbb{R}^{D}$ mirrors the $D=2$ case.
-Consider calculating the entanglement entropy between two codimension-one disjoint discs $A$ and $B$ in $\mathbb{R}^{D}$, corresponding to the mixed state $\rho_{AB}$ of the infinite system.
+Consider calculating the entanglement entropy between two codimension-one disjoint discs $A$ and $B$ in $\mathbb{R}^{D}$.
 Analogous to the two-dimensional scenario, removing a $D$-dimensional solid torus between $A$ and $B$ yields also a $D$-dimensional solid torus depicted in Figure 3 — precisely the SUBTRACTION procedure introduced in [3].
@@
 In section V, we show how RT formula [8, 9] naturally emerges in this framework and discuss arbitrary shape of entangling surfaces.
-In section VI, we find perfect agreement of our results with the holographic calculation.
+Section VI is the conclusion.
 II Entanglement entropy between two disjoint balls
-For a CFT in $D$-dimensional Euclidean spacetime with the metric $\mathrm{d}s_{0}^{2}=\mathrm{d}t_{\text{E}}^{2}+\mathrm{d}y^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2}$, we consider quantum fields living in a bounded region,
-:=$\mathcal{B}:=\left\{(t_{\text{E}},y,x)|\left(\sqrt{t_{\text{E}}^{2}+y^{2}}-\frac{l_{2}+l_{1}}{2}\right)^{2}+\sum_{i=1}^{D-2}x_{i}^{2}<\left(\frac{l_{2}-l_{1}}{2}\right)^{2}\right\},$ (1)
-with $l_{2}>l_{1}>0$.
-At the time slice $t_{\text{E}}=0$, $\mathcal{B}$ consists of two disjoint $(D-1)$-balls $A$ and $B$ with the same radius, as shown in Figure 3.
-The bounded region $\mathcal{B}$ is intuitively a solid torus $\mathbb{B}^{D-1}\times S^{1}$, where $\mathbb{B}^{D-1}$ is a $(D-1)$-dimensional ball.
-Up to a Weyl transformation $\mathrm{d}s_{0}^{2}\rightarrow\mathrm{d}s_{0}^{2}/r^{2}$, the spacetime metric becomes
-$\mathrm{d}s^{2}=\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\right)/r^{2},$ (2)
-in terms of coordinates transformations, $t_{\text{E}}=r\sin\theta$, $y=r\cos\theta$.
-In this parameterization, we can treat $\theta$ as imaginary time with a periodicity $2\pi$, and the CFT living in $\mathcal{B}$ can be thought as a thermal field theory with the inverse temperature $\beta=2\pi$.
-The partition function of the CFT can be formally written as a trace over the Hilbert space:
-$Z=\mathrm{Tr}\,e^{-2\pi H}.$ (3)
-The Hamiltonian operator $H$ conjugate to the imaginary time $\theta$ is defined by
+Figure 3: The Euclidean pure state density matrix $\rho=|\psi\rangle\langle\psi|$ for the solid torus CFTD.
+The lower half solid torus corresponds to $|\psi\rangle$.
+The upper half annulus corresponds to $\langle\psi|$.
+We calculate the entanglement entropy between two disjoint co-dimensional one balls AA and $B$ (colored regions) in the bounded region $\mathcal{B}$ (blue region).
+Let us consider the CFTD on a solid torus,
+$\mathcal{B}=\mathbb{B}^{D-1}\times S^{1}:=\left\{(t_{\text{E}},y,x)|\left(\sqrt{t_{\text{E}}^{2}+y^{2}}-\frac{l_{2}+l_{1}}{2}\right)^{2}+\sum_{i=1}^{D-2}x_{i}^{2}<\left(\frac{l_{2}-l_{1}}{2}\right)^{2}\right\},$ (1)
+where $\mathbb{B}^{D-1}$ is a (D1)(D-1)-dimensional ball and $l_{2}>l_{1}>0$.
+At the time slice $t_{\text{E}}=0$, $\mathcal{B}$ consists of two disjoint (D1)(D-1)-balls AA and $B$ with the same radius, as shown in Figure 3.
+The lower half solid torus corresponds to $|\psi\rangle$, and the upper half annulus corresponds to $\langle\psi|$.
+The density matrix of this pure state is $\rho=|\psi\rangle\langle\psi|$.
+Using the replica trick [1], the entanglement entropy $S_{\text{disj}}(A:B)$ between $A$ and $B$ can be defined through the Rényi entropy as
+$S^{(n)}(A:B)=\frac{1}{1-n}\log\mathrm{Tr}_{A}\rho_{A}^{n}=\frac{1}{1-n}\log\left[\frac{Z_{\mathcal{B}_{n}}}{Z^{n}}\right],$
+$S_{\text{disj}}(A:B)=\lim_{n\to 1}S^{(n)}(A:B),$ (2)
+where $\rho_{A}=\mathrm{Tr}_{B}\rho$ is the reduced density matrix.
+$Z$ is the partition function on the solid torus $\mathcal{B}$.
+$Z_{\mathcal{B}_{n}}$ is the partition function on the $n$-sheeted cover $\mathcal{B}_{n}$ made by cyclically gluing $n$ copies of $\mathcal{B}$ along $A$ (or equivalently $B$), see Figure 4 where the cut $A$ is a (D1)(D-1)-dimensional ball $\mathbb{B}^{D-1}$.
+Figure 4: The cut-and-glue procedure in the replica trick to construct $\mathcal{B}_{n}$.
+Each solid torus is cut along the (D1)(D-1)-dimensional ball $A$ and is glued with others cyclically.
+Red lines represent gluing operations.
+Note that the resulted manifold is also a solid torus with period $2n\pi$.
+Since $\mathcal{B}_{n}$ is still a solid torus with period $2n\pi$, calculating the $S_{\text{disj}}(A:B)$ in equation (2) is simplified to computing the partition function on a solid torus $\mathcal{B}=\mathbb{B}^{D-1}\times S^{1}$!
+A brute force approach to this calculation would involve solving the equations of motion (EoM) with either Dirichlet or Neumann boundary conditions which makes no difference since the vacuum entanglement is concerned.
+Instead, using the fact that a periodic system is equivalent to a thermal system, by noting the solid torus $\mathcal{B}$ is periodic, we adopt a symmetry based method: this approach is universal across all CFTs, applicable to free or strongly coupled systems.
+Applying the coordinate transformation $t_{\text{E}}=r\sin\theta$, $y=r\cos\theta$ to the flat metric $\mathrm{d}s_{0}^{2}=\mathrm{d}t_{\text{E}}^{2}+\mathrm{d}y^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2}$, we get a conformally flat metric
+$\mathrm{d}s_{0}^{2}=r^{2}\left[\mathrm{d}\theta^{2}+\Big(\mathrm{d}r^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\Big)/r^{2}\right].$
+Removing the conformal factor $r^{2}$ by the Weyl invariance, the spacetime metric becomes
+$\mathrm{d}s^{2}=\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\right)/r^{2}.$ (3)
+Since $\theta$ plays a role of the imaginary time with a periodicity $2\pi$, this CFT is equivalent to a thermal system with the inverse temperature $\beta=2\pi$.
+It is straightforward to write down the partition function of this thermal system
+$Z=\mathrm{Tr}\,e^{-2\pi H},$ (4)
+where the Hamiltonian operator $H$ is conjugate to the imaginary time $\theta$, defined by
 $H=\int_{\mathbb{B}^{D-1}}T^{\theta\theta}$
-where $T$ is the stress-energy tensor.
-Figure 3: Entangling regions $A$ and $B$ (colored regions) in the bounded region $\mathcal{B}$ (blue region).
-Starting point of our computations is the observation that the vacuum entanglement entropy between $A$ and $B$ can be computed by the replica trick [1].
-Such techniques allow us to write the Rényi entropy as
-$S^{(n)}=\frac{1}{1-n}\log\left[\frac{Z_{\mathcal{B}_{n}}}{Z^{n}}\right],$ (4)
-where $\mathcal{B}_{n}$ is made by cyclically gluing $n$ copies of $\mathcal{B}$ with a cut along $A$ (or equivalently $B$), and the metric of $\mathcal{B}_{n}$ might be given by
-$\mathrm{d}s_{(n)}^{2}=n^{2}\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\right)/r^{2}.$ (5)
-The partition function of CFT living on $\mathcal{B}_{n}$ is then
+where $T$ is the stress-energy tensor. By noting the periodicity of the $n$-sheeted cover $\mathcal{B}_{n}$ is $2n\pi$, we immediately get the partition function on $\mathcal{B}_{n}$,
 $Z_{\mathcal{B}_{n}}=\mathrm{Tr}\,e^{-2n\pi H_{(n)}},$ (5)
@@
 $\langle H_{(n)}\rangle=\frac{\langle H\rangle}{n^{2}}.$ (7)
-Then, the Rényi entropy becomes
-$S^{(n)}=\frac{1}{1-n}\log\left[\frac{e^{-2\pi\langle H\rangle/n}}{e^{-2n\pi\langle H\rangle}}\right]=-2\pi\left(1+\frac{1}{n}\right)\langle H\rangle.$ (9)
-The entropy $S(A:B)$ of entanglement between two disjoint balls $A,B$ therefore reads
-$S(A:B)=S^{(1)}=-4\pi\langle H\rangle.$ (10)
+Then, the Rényi entropy in equation (2) becomes
+$S^{(n)}(A:B)=\frac{1}{1-n}\log\left[\frac{e^{-2\pi\langle H\rangle/n}}{e^{-2n\pi\langle H\rangle}}\right]=-2\pi\left(1+\frac{1}{n}\right)\langle H\rangle.$ (8)
+The entropy $S_{\text{disj}}(A:B)$ of entanglement between two disjoint balls $A,B$ therefore reads
+$S_{\text{disj}}(A:B)=S^{(1)}(A:B)=-4\pi\langle H\rangle.$ (9)
 We stress that the Hamiltonian $H$ is conjugate to $\theta$ but not to $t_{\text{E}}$, so it is actually dimensionless.
-The entanglement entropy $S(A:B)$ simply depends on the ground state energy $\langle H\rangle$ localized in a spatial ball $\mathbb{B}^{D-1}$:
+The entanglement entropy $S_{\text{disj}}(A:B)$ simply depends on the ground state energy $\langle H\rangle$ localized in a spatial ball $\mathbb{B}^{D-1}$:
 $\langle H\rangle=\int_{\mathbb{B}^{D-1}}\mathcal{E}_{\text{vac}},$ (10)
-where $\mathcal{E}_{\text{vac}}$ is the Casimir energy density that is usually a negative constant in our discussions. Consequently, the entanglement entropy $S(A:B)$ is given by
-$S(A:B)$ $=$ $-4\pi\mathcal{E}_{\text{vac}}\mathrm{Vol}(\mathbb{B}^{D-1}),$ (12)
+where $\mathcal{E}_{\text{vac}}$ is the Casimir energy density that is usually a negative constant in our discussions.
+The integration is performed with the metric of equation (3).
+Consequently, the entanglement entropy $S_{\text{disj}}(A:B)$ is given by
+$S_{\text{disj}}(A:B)$ $=$ $-4\pi\mathcal{E}_{\text{vac}}\mathrm{Vol}(\mathbb{B}^{D-1}),$ (11)
 $=$ $-4\mathcal{E}_{\text{vac}}\pi^{\frac{D}{2}}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{l_{2}}{l_{1}}-1\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;1-\frac{l_{2}}{l_{1}}\right).$
-with the standard hypergeometric function $\,{}_{2}F_{1}$. The parameters $l_{1}$ and $l_{2}$ are the inner and outer radii of the solid torus, as indicated in the left panel of Figure 5.
-Figure 4: Entangling regions (colored regions) are bounded by ($D-2$)-spheres (black circles), respectively.
-Left panel (the cavity configuration): A ($D-1$)-ball living in a spheric cavity is entangled with the region outside the cavity.
-Right panel (the juxtaposed configuration): Two disjoint ($D-1$)-balls with different radii are entangled.
+with the standard hypergeometric function $_{2}F_{1}$. The parameters $l_{1}$ and $l_{2}$ are the inner and outer radii of the solid torus.
+Figure 5: Entangling regions (colored regions) are bounded by (D2D-2)-spheres (black circles), respectively.
+Left panel (the cavity configuration): A (D1D-1)-ball living in a spheric cavity is entangled with the region outside the cavity.
+Right panel (the juxtaposed configuration): Two disjoint (D1D-1)-balls with different radii are entangled.
 Generally, there exist two different configurations of our interest, namely, the cavity configuration and the juxtaposed configuration illustrated in Figure 5.
-Each configuration is uniquely fixed by boundaries of $A,B$, i.e., two disjoint ($D-2$)-spheres.
-Every ($D-1$)-ball is bounded by a ($D-2$)-sphere $\Sigma(\vec{x},r)$ located at $\vec{x}\in\mathbb{R}^{D-1}$ with the radius $r$, respectively.
+Each configuration is uniquely fixed by boundaries of $A,B$, i.e., two disjoint (D2D-2)-spheres.
+Every (D1D-1)-ball is bounded by a (D2D-2)-sphere $\Sigma(\vec{x},r)$ located at $\vec{x}\in\mathbb{R}^{D-1}$ with the radius $r$, respectively.
 An elegant substitute for conformality is the inversive product $\varrho$ of two spheres $\Sigma(\vec{x},r)$ and $\Sigma(\vec{x}^{\prime},r^{\prime})$ [10]
 $\varrho=\left|\frac{r^{2}+r^{\prime 2}-|\vec{x}-\vec{x}^{\prime}|^{2}}{2rr^{\prime}}\right|.$ (12)
-In the case that two disjoint ($D-1$)-balls $A,B$ share the same radius, we have
+In the case that two disjoint (D1D-1)-balls $A,B$ share the same radius, we have
 $\frac{l_{2}}{l_{1}}=\frac{\sqrt{\varrho+1}+\sqrt{2}}{\sqrt{\varrho+1}-\sqrt{2}}.$ (13)
@@
 Consequently, any two configurations sharing the same $\varrho$ must be conformally equivalent.
-To compute the entanglement entropy $S(A:B)$ for any two disjoint spheres in general configurations, it suffices to evaluate the inversive product $\varrho$ of the spheres.
-Thus, we conclude that the entanglement entropy between any two disjoint spheres is given by [^2] For the time dependent covariant scenario, simply replace $\vec{x}$ with $x^{\mu}$.
-$S=-4\mathcal{E}_{\text{vac}}\pi^{\frac{D}{2}}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{2\sqrt{2}}{\sqrt{\varrho+1}-\sqrt{2}}\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;\frac{2\sqrt{2}}{\sqrt{2}-\sqrt{\varrho+1}}\right).$ (15)
+To compute the entanglement entropy $S_{\text{disj}}(A:B)$ for any two disjoint spheres in general configurations, it suffices to evaluate the inversive product $\varrho$ of the spheres.
+Thus, we conclude that the entanglement entropy between any two disjoint spheres is given by
+$S_{\text{disj}}(A:B)=-4\mathcal{E}_{\text{vac}}\pi^{\frac{D}{2}}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{2\sqrt{2}}{\sqrt{\varrho+1}-\sqrt{2}}\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;\frac{2\sqrt{2}}{\sqrt{2}-\sqrt{\varrho+1}}\right).$ (14)
+The time dependent covariant scenario is obtained by simply replacing $\vec{x}$ with $x^{\mu}$.
 III Emergent area law in various dimensions
 In this section, we take the adjacent limit of the general finite entanglement entropy (14) of disjoint balls to give the usual divergent entanglement entropies of adjacent balls.
-Figure 5: Left panel: The time zero slice of $\mathcal{B}$ in $D$ dimensions, where $A$ and $B$ are two disjoint ($D-1$)-balls with the same radius (colored regions) living at $t_{\text{E}}=0$.
-Right panel: The ($D-1$)-ball $A$ with the radius $R-\epsilon$ is in a spheric cavity with the radius $R+2\epsilon$, and the region $B$ is outside the cavity.
+Figure 6: Left panel: The time zero slice of $\mathcal{B}$ in $D$ dimensions, where $A$ and $B$ are two disjoint (D1D-1)-balls with the same radius (colored regions) living at $t_{\text{E}}=0$.
+Right panel: The (D1D-1)-ball AA with the radius $R-\epsilon$ is in a spheric cavity with the radius $R+2\epsilon$, and the region $B$ is outside the cavity.
 Here, $\epsilon$ plays a role of the UV cutoff.
-We first consider the entanglement entropy $S(A:B)$ for two disjoint ($D-1$)-balls $A,B$ with the same radius in the juxtaposed configuration, as shown in Figure 5.
-It would be convenient to write down $S(A:B)$ in various dimensions:
+We first consider the entanglement entropy $S_{\mathrm{disj}}(A:B)$ for two disjoint (D1D-1)-balls $A,B$ with the same radius in the juxtaposed configuration, as shown in Figure 6.
+It would be convenient to write down $S_{\text{disj}}(A:B)$ in various dimensions:
 $D=2:$ $4\pi\mathcal{E}_{2}\log\frac{l_{2}}{l_{1}},$
@@
 where we denote $\mathcal{E}_{\text{vac}}=-\mathcal{E}_{D}<0$ and remember that $\mathcal{E}_{D}$ is actually different in various dimensions.
-Then, we return to the general $D$-dimensional cavity configuration depicted in the right panel of Figure 5, where $A$ is almost adjacent to $B$, more precisely implying that
+Then, we return to the general $D$-dimensional cavity configuration depicted in the right panel of Figure 6, where $A$ is almost adjacent to $B$, more precisely
 $\vec{x}=\vec{x}^{\prime},\quad r=R+\epsilon,\quad r^{\prime}=R-\epsilon.$ (16)
@@
 $\varrho=\frac{R^{2}+\epsilon^{2}}{R^{2}-\epsilon^{2}}.$ (17)
-Plugging $\varrho$ into Equation (16), we obtain the corresponding entanglement entropy in various dimensions:
+Plugging $\varrho$ into the general expression (14), we obtain $S_{\text{adj}}(A:B)$ in various dimensions:
 $D=2:$ $8\pi\mathcal{E}_{2}\log\frac{2R}{\epsilon}+\mathcal{O}(\epsilon),$
@@
 $D=5:$ $\frac{8\pi^{3}\mathcal{E}_{5}}{3}\left(\frac{R^{3}}{\epsilon^{3}}-\frac{3R}{\epsilon}+2\right),$
-$D=6:$ $\frac{\pi^{3}\mathcal{E}_{6}}{3}\left(\frac{8R^{4}}{\epsilon^{4}}-\frac{24R^{2}}{\epsilon^{2}}+9+12\log\frac{2R}{\epsilon}\right)+\mathcal{O}\left(\epsilon\right).$ (19)
+$D=6:$ $\frac{\pi^{3}\mathcal{E}_{6}}{3}\left(\frac{8R^{4}}{\epsilon^{4}}-\frac{24R^{2}}{\epsilon^{2}}+9+12\log\frac{2R}{\epsilon}\right)+\mathcal{O}\left(\epsilon\right),$ (18)
+which perfectly agrees with holographic predictions given in [8].
 In $D=2$, the result is the well-known $\log$-law for entanglement entropy.
@@
 IV Examples in CFT2 and CFT4
-We provide two consistency checks in this section.
+We provide two checks in this section.
 Consider a CFT4 living in the 4-dimensional conformally flat spacetime with the metric $\mathrm{d}s^{2}=\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{2}\mathrm{d}x_{i}^{2}\right)/r^{2}$, where the vacuum expectation value of the stress tensor has been explicitly calculated in [14],
@@
 where $a$ and $c$ are different coefficients of the trace anomaly [^2] In our notations, the trace anomaly is $\left\langle T_{\;\mu}^{\mu}\right\rangle=\frac{1}{(4\pi)^{2}}\left[c(C_{\mu\nu\rho\sigma}^{2}+\frac{2}{3}\nabla^{2}\mathcal{R})-a\mathcal{G}\right]$, with the Weyl tensor $C_{\mu\nu\rho\sigma}$, the Ricci scalar $\mathcal{R}$ and the topological Euler density $\mathcal{G}.$.
-Plugging Equation (20) and $D=4$ into Equation (12), the entanglement entropy between $A$ and $B$ is given by
-$S(A:B)=\frac{3a+c}{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-\frac{3a+c}{2}\log\frac{l_{2}}{l_{1}}.$ (21)
-Note that $A$, $B$ are two disjoint 3-balls with the same radius. In the 4-dimensional $\mathcal{N}=4$ $SU(N)$ super Yang-Mills gauge theory, the anomaly coefficients are [15, 16]
+In the 4-dimensional $\mathcal{N}=4$ $SU(N)$ super Yang-Mills gauge theory, the anomaly coefficients are [15, 16]
 $a=\frac{N^{2}}{4},\quad c=\frac{N^{2}}{4}.$ (20)
-Hence
-$S(A:B)=\frac{N^{2}}{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-\frac{N^{2}}{2}\log\frac{l_{2}}{l_{1}}.$ (23)
-Traditionally, that $A$ and $B$ are two adjacent and complementary regions is of our main interest, as depicted in the right panel of Figure 5.
-This case is conformally equivalent to two disjoint 3-balls with the same radius, due to the fact that any global conformal transformations are sphere-preserving [10], which has been analyzed in a more general way in Section III.
-Consequently, in Figure 5, the left panel and the right panel are conformally equivalent by setting that
-$\frac{l_{2}}{l_{1}}\simeq\left(\frac{2R}{\epsilon}\right)^{2}-2+\mathcal{O}(\epsilon),$ (24)
-if we take the adjacent limit $\epsilon\rightarrow 0$. Such a setup finally gives us the area law of entanglement entropy:
-$S(A:A^{c})\simeq N^{2}\left[\left(\frac{R}{\epsilon}\right)^{2}-\log\frac{2R}{\epsilon}-2\right]+\mathcal{O}(\epsilon).$ (25)
-This is actually entanglement entropy of a ball with the radius $R$, which precisely matches the holographic prediction given in [8].
+Hence from equation (15), for the juxtaposed configuration,
+$S_{\text{disj}}(A:B)=\frac{N^{2}}{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-\frac{N^{2}}{2}\log\frac{l_{2}}{l_{1}}.$ (21)
+For the right panel of Figure 6, from equation (18), the adjacent entanglement entropy $S_{\text{adj}}(A:A^{c})$ of a 33-ball in $\mathbb{R}^{3}$ is
+$S_{\text{adj}}(A:A^{c})\simeq N^{2}\left[\left(\frac{R}{\epsilon}\right)^{2}-\log\frac{2R}{\epsilon}-2\right]+\mathcal{O}(\epsilon),$ (22)
+which precisely matches the holographic prediction given in [8].
 As $D=2$, the metric reduces to $\mathrm{d}s^{2}=\mathrm{d}\theta^{2}+\mathrm{d}r^{2}/r^{2}$, which is just a flat metric on the Euclidean cylinder.
@@
 Its adjacent limit gives the entanglement entropy of a single interval with the length $2R$,
-$S(A:A^{c})=\frac{c}{3}\log\frac{2R}{\epsilon}+\mathcal{O}(\epsilon).$
+$S_{\text{adj}}(A:A^{c})=\frac{c}{3}\log\frac{2R}{\epsilon}+\mathcal{O}(\epsilon).$
 V Shape independence and Ryu-Takayanagi formula
 We present two important indications or extensions in this section.
-Shape independence
-Even though we only study the spheric entangling surface in this paper, our approach is clearly independent of the specific shape of entangling regions $A$ and $B$.
-In other words, one can replace $A$ and $B$ with generally shaped regions, denoted as $\mathbb{X}$.
-The relevant entanglement entropy would then be
-$S(A:B)$ $=$ $-4\pi\mathcal{E}_{\text{vac}}\mathrm{Vol}(\mathbb{X}).$ (28)
-However, it should be noted that to apply this formula directly, $A$ and $B$ must be reflection-symmetric with respect to each other, as only under this condition the thermal partition function can be formulated.
-To calculate the entanglement entropy between two similar, disjoint regions (same shape, different sizes) with arbitrary positions, one needs to map the configuration to the reflection-symmetric one with some conformal map.
-For any shape of the entangling region, there should exist a set of inversive products analogous to (13).
-For adjacent scenarios, it is easy to understand that, by choosing some simple setup and then taking adjacent limit, the usual divergent entanglement entropy for arbitrarily shaped region can be obtained straightforwardly.
-Manifestation of Ryu-Takayanagi formula
-Since any pair of similar, disjoint regions falls into the reflection-symmetric case, the entanglement entropy is proportional to the area $\mathrm{Vol}(\mathbb{X})$ defined within the hyperbolic metric $\mathrm{d}s^{2}=(\mathrm{d}r^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2})/r^{2}.$ Note that the subregion $\mathbb{X}$ could always be understood as a ($D-1$)-dimensional minimal surface embedded in higher-dimensional AdSD+1 space.
+V.1 Shape independence
+Although we only study the spherical entangling surface in this paper, our approach is clearly independent of the specific shape of entangling regions $A$ and $B$ in the reflection-symmetric juxtaposed configuration.
+One simply replaces $A$ and $B$ with a generally shaped region, denoted as $\mathbb{X}$.
+The entanglement entropy would then be
+$S_{\text{disj}}(A:B)$ $=$ $-8\pi\mathcal{E}_{\text{vac}}\mathrm{Vol}(\mathbb{X}),$ (25)
+where $\mathrm{Vol}(\mathbb{X})$ is easily calculated with the conformal metric (3).
+A particular interesting object is the divergent adjacent entanglement entropy $S_{\text{adj}}(\mathbb{Y}:\mathbb{Y}^{c})$ for an arbitrarily shaped region $\mathbb{Y}$ in $\mathbb{R}^{D}$.
+It can be calculated as follows:
+  • We start with $S_{\text{disj}}(\mathbb{Y}_{1}:\mathbb{Y}_{2})$ in the cavity configuration. $\mathbb{Y}_{1}$ and $\mathbb{Y}_{2}$ are nested, centered at the same point, different in scale.
+  • To calculate $S_{\text{disj}}(\mathbb{Y}_{1}:\mathbb{Y}_{2})$, we need to map it to the juxtaposed configuration.
+  • Typically we end up with a non-reflection-symmetric juxtaposed configuration.
+  • Since the thermal partition function usually cannot be formulated for generic non-reflection-symmetric juxtaposed configurations, analytic expression of $S_{\text{disj}}(\mathbb{Y}_{1}:\mathbb{Y}_{2})$ may not be available.
+In this case, we can compute the partition function on the non-reflection-symmetric torus numerically.
+  • Finally, take the adjacent limit,
+    $S_{\text{adj}}(\mathbb{Y}:\mathbb{Y}^{c})=\lim_{\mathbb{Y}_{1}\to\mathbb{Y}_{2}}S_{\text{disj}}(\mathbb{Y}_{1}:\mathbb{Y}_{2}).$
+V.2 Manifestation of Ryu-Takayanagi formula
+Since any pair of similar, disjoint regions falls into the reflection-symmetric case, the entanglement entropy is proportional to the volume $\mathrm{Vol}(\mathbb{X})$ defined within the hyperbolic metric $\mathrm{d}s^{2}=(\mathrm{d}r^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2})/r^{2}.$ Note that the subregion $\mathbb{X}$ could always be understood as a (D1D-1)-dimensional minimal surface embedded in AdSD+1 space, up to a conformal transformation for asymmetric configurations.
 Therefore, the RT formula saying that the entanglement entropy of CFTD is proportional to the area of the codimension-2 minimal surface in the AdSD+1 spacetime has been manifested in current method.
-VI Holographic entanglement entropy
-Figure 6: At the time slice $t_{\text{E}}=0$, two subregions $A,B$ are living on the AdS boundary $z=0$, while the minimal cross-section $E_{W}$ (gray region) of the entanglement wedge (the region bounded by the blue tube) of $A$ and $B$ is in the AdS bulk.
-It would be nice to explore the gravity dual of finite entanglement entropy in the context of AdS/CFT correspondence [18, 19, 20].
-Recall that a CFT in $D$-dimensional Euclidean spacetime with the metric $\mathrm{d}s_{0}^{2}=\mathrm{d}t_{\text{E}}^{2}+\mathrm{d}y^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2}$ is dual to a ($D+1$)-dimensional Euclidean AdS spacetime with the metric
-$\mathrm{d}s^{2}=\frac{\ell_{\text{AdS}}^{2}}{z^{2}}\left(\mathrm{d}z^{2}+\mathrm{d}t_{\text{E}}^{2}+\mathrm{d}y^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\right).$ (29)
-The entanglement entropy between two disjoint $(D-1)$-balls $A$ and $B$ with the same radius is dual to a ball-shaped region at the time slice $t_{\text{E}}=0$ of the AdSD+1 bulk, i.e., the entanglement wedge cross-section,
-$E_{W}:=\left\{(x,z)|\left(z-\frac{l_{2}+l_{1}}{2}\right)^{2}+\sum_{i=1}^{D-2}x_{i}^{2}<\left(\frac{l_{2}-l_{1}}{2}\right)^{2}\right\}.$ (30)
-The area of the entanglement wedge cross-section is given by
-$\mathsf{A}=\ell_{\text{AdS}}^{D-1}\pi^{\frac{D}{2}-1}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{l_{2}}{l_{1}}-1\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;1-\frac{l_{2}}{l_{1}}\right).$ (31)
-According to the RT formula, the corresponding entanglement entropy is
-$S(A:B)=\frac{\mathsf{A}}{4G_{N}^{(D+1)}}=\frac{\ell_{\text{AdS}}^{D-1}}{4G_{N}^{(D+1)}}\pi^{\frac{D}{2}-1}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{l_{2}}{l_{1}}-1\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;1-\frac{l_{2}}{l_{1}}\right),$ (32)
-with the ($D+1$)-dimensional gravitational constant $G_{N}^{(D+1)}$. This coincides with our field-theoretic result (15). For instance, in $D=4$, one has
-$S(A:B)$ $=$ $\frac{3a+c}{4}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}\right)-\frac{3a+c}{2}\log\frac{l_{2}}{l_{1}}$ (33)
-$=$ $\frac{\pi\ell_{\text{AdS}}^{3}}{8G_{N}^{(5)}}\left(\frac{l_{2}}{l_{1}}-\frac{l_{1}}{l_{2}}-2\log\frac{l_{2}}{l_{1}}\right).$ (34)
-with two central charges $a=c=\pi\ell_{\text{AdS}}^{3}/8G_{N}^{(5)}$ [21, 22].
-It is straightforward to extend the entanglement wedge cross-section to the general case, such as the cavity configuration we have addressed in the CFT side.
-This is because the AdS bulk and the dual CFT always share the same conformal symmetry.
-The entanglement wedge cross-section in the general configuration can be conformally mapped onto the reflection-symmetric case studied above.
 When taking the adjacent limit, one immediately reproduces the original RT formula for the adjacent scenario with divergent entanglement entropy.
-VII Conclusion
+It is very interesting to note that, the disjoint entanglement entropy $S_{\text{disj}}(A:B)$ is proportional to the scaled volume of the entangling region., rather than the area of the entangling surface.
+As taking the adjacent limit, the contribution from the entangling surface dominates and the area law emerges.
+VI Conclusion
 In summary, we introduced a general framework to compute the entanglement entropy in CFTD for $D\geq 2$.
@@
 The RT formula emerges naturally within this formalism.
+In [18], some parts of finite contributions to the adjacent entanglement entropy were extracted for $D>2$ under specific limits, including: (i) entropy between the interior and exterior of a spatial domain for a massive scalar field; (ii) numerically computed entropy for an interval in CFT.
+Extending our framework to massive QFT is a critical research avenue.
+Additionally, [19] established the universal logarithmic term of adjacent entanglement entropy in even-dimensional CFT.
+A particularly intriguing direction is to recover the hyperbolic mapping introduced in that work from the adjacent limits of our results.
 In CFT2, the divergent entanglement entropy between two adjacent and complementary subsystems has been extensively studied.

### v3

@@
 The upper half annulus corresponds to $\langle\psi|$.
-We calculate the entanglement entropy between two disjoint co-dimensional one balls AA and $B$ (colored regions) in the bounded region $\mathcal{B}$ (blue region).
+We calculate the entanglement entropy between two disjoint co-dimensional one balls $A$ and $B$ (colored regions) in the bounded region $\mathcal{B}$ (blue region).
 Let us consider the CFTD on a solid torus,
@@
 $S_{\text{adj}}(A:A^{c})=\frac{c}{3}\log\frac{2R}{\epsilon}+\mathcal{O}(\epsilon).$
-V Shape independence and Ryu-Takayanagi formula
+V Ryu-Takayanagi formula and Shape independence
 We present two important indications or extensions in this section.
-V.1 Shape independence
+V.1 Manifestation of Ryu-Takayanagi formula
+Our derivation has already manifested the RT formula.
+From the first line of equations (11), for a symmetric solid torus (reflection-symmetric juxtaposed configuration) as illustrated by the left panel of Figure 6, the entanglement entropy between balls $A$ and $B$ is the volume Vol(A)(A) (or equivalently Vol(B)(B)) calculated in the hyperbolic metric $\mathrm{d}s^{2}=(\mathrm{d}r^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2})/r^{2}.$ This hyperbolic volume Vol(A)(A) could always be understood as a (D1D-1)-dimensional minimal surface embedded in AdSD+1 space, which is precisely the entanglement wedge cross-section (EWCS), i.e RT surface, for this symmetric configuration.
+As for a generic disjoint configurations in Figure 5, we simply make a conformal transformation to get the corresponding RT surface (EWCS).
+This actually provides a simple way to compute the minimal surface for a particular configuration.
+It is of interest to verify the result in a geometric picture.
+When taking the adjacent limit, one immediately reproduces the original RT formula for the adjacent scenario with divergent entanglement entropy.
+It is very interesting to note that, the disjoint entanglement entropy $S_{\text{disj}}(A:B)$ is proportional to the hyperbolic volume of the entangling region., rather than the area of the entangling surface.
+As taking the adjacent limit, the contribution from the entangling surface dominates and the area law emerges.
+V.2 Shape independence
 Although we only study the spherical entangling surface in this paper, our approach is clearly independent of the specific shape of entangling regions $A$ and $B$ in the reflection-symmetric juxtaposed configuration.
@@
     $S_{\text{adj}}(\mathbb{Y}:\mathbb{Y}^{c})=\lim_{\mathbb{Y}_{1}\to\mathbb{Y}_{2}}S_{\text{disj}}(\mathbb{Y}_{1}:\mathbb{Y}_{2}).$
-V.2 Manifestation of Ryu-Takayanagi formula
-Since any pair of similar, disjoint regions falls into the reflection-symmetric case, the entanglement entropy is proportional to the volume $\mathrm{Vol}(\mathbb{X})$ defined within the hyperbolic metric $\mathrm{d}s^{2}=(\mathrm{d}r^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2})/r^{2}.$ Note that the subregion $\mathbb{X}$ could always be understood as a (D1D-1)-dimensional minimal surface embedded in AdSD+1 space, up to a conformal transformation for asymmetric configurations.
-Therefore, the RT formula saying that the entanglement entropy of CFTD is proportional to the area of the codimension-2 minimal surface in the AdSD+1 spacetime has been manifested in current method.
-When taking the adjacent limit, one immediately reproduces the original RT formula for the adjacent scenario with divergent entanglement entropy.
-It is very interesting to note that, the disjoint entanglement entropy $S_{\text{disj}}(A:B)$ is proportional to the scaled volume of the entangling region., rather than the area of the entangling surface.
-As taking the adjacent limit, the contribution from the entangling surface dominates and the area law emerges.
 VI Conclusion